OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

iYuko (iyuko):

y2 - y1 ---------- = slope x2 - x1

iYuko (iyuko):

-4= x1 8= y1

iYuko (iyuko):

-3=x2 -6=y2

OpenStudy (anonymous):

so would the answer be -14